The force stretching the D string on a certain guitar is 166N. The strings linear mass density is 0.006kg/m. What is the speed of the waves on the string? In m/s


1
Expert's answer
2021-10-06T08:13:08-0400

We can find the speed of the waves on the string as follows:


"v=\sqrt{\dfrac{T}{\mu}}=\sqrt{\dfrac{166\ N}{0.006\ \dfrac{kg}{m}}}=166\ \dfrac{m}{s}."